what is the wavelength of light travelling with frequeny of 55hz ?

Grade:10

1 Answers

Arun
5 years ago
Dear Vikas
 
v = f * lambda
 
3 * 10^8 = 55 * lambda
 
lambda = 300 * 10^6 /55
 
 = 5.45 * 10^6 metres
 
Hope it helps
 
Regards
Arun (askIITians forum expert)